# Dynamical Similarity in Multisymplectic Field Theory

### v3

@@ II.3 Hamiltonian Field Theory
Having discussed the Lagrangian formulation of multisymplectic field theory, we would now like to consider the Hamiltonian counterpart. The construction of a covariant Hamiltonian formalism is less straightforward, since there is no uniquely-defined multiphase space [46, 47, 48]. In what follows, we shall adopt the simplest and most ubiquitous framework found throughout the literature, and construct the Hamiltonian formalism using the so-called extended and restricted multimomentum bundles [49, 50].
Let us denote by $\wedge^{m}_{2}\,T^{*}E$ the space of differential $m$-forms on $E$ that vanish when acting upon two $\pi$-vertical vector fields. We shall refer to $\wedge^{m}_{2}\,T^{*}E$ as the extended multimomentum bundle; its local coordinates adapted to the underlying bundle structure are $(x^{\mu},y^{a},p^{\mu}_{a},p)$, and there exist two natural submersions
$\mu:\wedge^{m}_{2}\,T^{*}E\rightarrow E\quad\quad\textrm{and}\quad\quad\bar{\mu}=\pi\circ\mu:\wedge^{m}_{2}\,T^{*}E\rightarrow M$
The extended multimomentum bundle $\wedge^{m}_{2}\,T^{*}E$ admits a differential $m$-form $\Theta$, known as the Liouville form, which, in local coordinates, is expressed as
$\Theta=p^{\mu}_{a}\,dy^{a}\wedge d^{m-1}x_{\mu}+p\,d^{m}x$ (10)
This Liouville $m$-form then allows us to define a multisymplectic form $\Omega=-\,d\Theta\in\Omega^{m+1}(\wedge^{m}_{2}\,T^{*}E)$; however, the extended multimomentum bundle $\wedge^{m}_{2}\,T^{*}E$ is not the arena in which multisymplectic Hamiltonian dynamics takes place. In order to define this manifold, we first consider the bundle whose sections are $\pi$-semibasic $m$-forms on $E$, which we denote by $\pi^{*}\hskip-2.56073pt\wedge^{m}\hskip-1.42262ptT^{*}M$, and introduce the quotient
$J^{1}E^{*}:=\frac{\wedge^{m}_{2}\,T^{*}E}{\pi^{*}\hskip-2.56073pt\wedge^{m}\hskip-1.42262ptT^{*}M}$
This space is made into a bundle, which we call the restricted multimomentum bundle, upon defining the following projection maps
$\sigma:J^{1}E^{*}\rightarrow E\quad\quad\quad\quad\bar{\sigma}=\pi\circ\sigma:J^{1}E^{*}\rightarrow M$
The restricted multimomentum bundle will serve as our covariant multiphase space for Hamiltonian dynamics. Local coordinates on $J^{1}E^{*}$, adapted to the bundle structure of $E$, are $(x^{\mu},y^{a},p^{\mu}_{a})$. Additionally, we have a naturally-defined map $\widetilde{\mathcal{FL}}:J^{1}E\rightarrow\wedge^{m}_{2}T^{*}E$, known as the extended Legendre transform; this is then restricted to the map $\mathcal{FL}:=\mu\circ\widetilde{\mathcal{FL}}:J^{1}E\rightarrow J^{1}E^{*}$, which acts on local bundle coordinates according to
@@
In complete analogy to the standard symplectic case, we shall call our Lagrangian system $(J^{1}E,\Omega_{L})$ regular if $\mathcal{FL}$ is a local diffeomorphism, hyperregular if this diffeomorphism is global, and singular if $\mathcal{FL}$ is not a diffeomorphism [51, 52, 53]. Additionally, we shall further specify a singular Lagrangian to be almost-regular [54], if the following three conditions are satisfied:
  • $\mathcal{I}=\mathcal{FL}(J^{1}E)$ is a closed submanifold of $J^{1}E^{*}$
  • The map $\mathcal{FL}$ is a submersion onto its image
  • The fibres $\mathcal{FL}^{-1}(p)$ are connected submanifolds of $J^{1}E$ for all points $p\in\mathcal{I}$
The regular and hyperregular cases are treated in an identical manner, with the exception that the regular case requires us to work locally on $J^{1}E$, rather than globally. The almost-regular case, on the other hand, must be addressed separately, through use of a suitable constraint algorithm [42]; further comments will be made on the treatment of singular Lagrangians in due course, as they are particularly relevant to the discussion of gauge theories within the multisymplectic framework [55, 56].
Pulling back the Liouville forms $\Theta$ and $\Omega$ along a Hamiltonian section $h:J^{1}E^{*}\rightarrow\wedge^{m}_{2}\,T^{*}E$, we obtain corresponding $m$ and $(m+1)$-forms on $J^{1}E^{*}$
$\Theta_{h}=h^{*}\Theta\;\in\;\Omega^{m}(J^{1}E^{*})\quad\quad\quad\textrm{and}\quad\quad\quad\Omega_{h}=-\,d\Theta_{h}\;\in\;\Omega^{m+1}(J^{1}E^{*})$ (12)
We refer to these objects as the Hamilton-Cartan $m$ and $(m+1)$-forms, and further define the pair $(J^{1}E^{*},\Omega_{h})$ to be a multisymplectic Hamiltonian system. The Hamiltonian function $H:\mathcal{U}\subset J^{1}E^{*}\rightarrow\mathbb{R}$ is then defined in terms of the section $h$ according to the following local coordinate expression
$h(x^{\mu},y^{a},p^{\mu}_{a})=\left(x^{\mu},y^{a},p^{\mu}_{a},p=-\,H(x^{\nu},y^{b},p^{\nu}_{b})\right)$ (13)
With this definition of the Hamiltonian function, the Hamilton-Cartan forms $\Theta_{h}$ and $\Omega_{h}$ are expressed locally as
$\begin{split}\Theta_{h}&=p^{\mu}_{a}\,dy^{a}\wedge d^{m-1}x_{\mu}-H\,d^{m}x\\ \vskip 0.56905pt\Omega_{h}=-\,&dp^{\mu}_{a}\wedge dy^{a}\wedge d^{m-1}x_{\mu}+dH\wedge d^{m}x\end{split}$ (14)
From $\Omega_{h}$, we shall also define the following 2-forms
$\omega_{h}^{\mu}:=-\,\iota_{\partial_{m}}\cdots\,\iota_{\partial_{1}}(\Omega_{h}\wedge dx^{\mu})=dy^{a}\wedge\,dp^{\mu}_{a}$ (15)
which will be used extensively in section (IV.2), when constructing vector fields corresponding to dynamical similarities.
In a similar spirit to the results of section (II.2), the Hamiltonian field equations, like their Lagrangian counterpart, may also be deduced from a variational principle. This so-called Hamilton-Jacobi principle seeks a critical section $\psi\in\Gamma(M,J^{1}E^{*})$, which, in this context is defined as
$\psi\;\,\textrm{critical}\quad\quad\Longleftrightarrow\quad\quad\frac{d}{dt}\biggr|_{t=0}\,\int_{M}\psi_{t}^{*}\Theta_{h}=0$
In a similar manner to the corresponding Lagrangian principle, the variations $\psi_{t}=f_{t}\circ\psi$ are constructed using a one-parameter family of transformations $\{f_{t}\}$ which generate any compactly-supported $\bar{\sigma}$-vertical vector field on $J^{1}E^{*}$.
The equivalent geometrical formulation of the Hamilton-Jacobi principle seeks an $m$-dimensional, $\bar{\sigma}$-transverse, integrable distribution $\mathcal{D}$, whose integral manifolds are critical sections of this variational problem. As in the Lagrangian case, such distributions are associated with classes of locally decomposable, $\bar{\sigma}$-transverse multivector fields $\{\boldsymbol{X}\}$, which may be expressed locally as
$\boldsymbol{X}=\bigwedge\limits_{\mu=1}^{m}\,f\left(\frac{\partial}{\partial x^{\mu}}+F^{a}_{\mu}\,\frac{\partial}{\partial y^{a}}+G^{\nu}_{a\mu}\,\frac{\partial}{\partial p^{\nu}_{a}}\right)$ (16)
for some non-vanishing function $f\in C^{\infty}(J^{1}E^{*})$. Continuing to draw parallels between the Lagrangian and Hamiltonian cases, if $\psi\in\Gamma(M,J^{1}E^{*})$ is a critical section of $\bar{\sigma}$, then
$\psi^{*}\,\iota_{\scriptscriptstyle X}\Omega_{h}=0\quad\quad\quad\quad\textrm{for all}\;X\in\mathfrak{X}^{\infty}(J^{1}E^{*})$ (17)
Additionally, if we express $\psi$ locally as $\psi(x)=(x^{\mu},\psi^{a}(x),\psi^{\mu}_{a}(x))$, since (17) holds for all vector fields $X\in\mathfrak{X}^{\infty}(J^{1}E^{*})$, using the local expression (14) for the Hamilton-Cartan form $\Omega_{h}$, we obtain the following Hamilton De-Donder Weyl equations
@@
Finally, $\psi$ is an integral section of a class of $\bar{\sigma}$-transverse integrable multivector fields $\{\boldsymbol{X}_{h}\}$, each of which satisfies
$\iota_{\scriptscriptstyle\boldsymbol{X}_{h}}\Omega_{h}=0$ (19)
We conclude this discussion describing the relationship between the critical sections of each of the two variational principles introduced thus far. Suppose that $\phi\in\Gamma(M,E)$ is a critical section for the (hyperregular) Lagrangian system $(J^{1}E,\Omega_{L})$. The object $\psi=\mathcal{FL}\circ j^{1}\phi\in\Gamma(M,J^{1}E^{*})$ then satisfies the Hamilton-Jacobi variational problem. Conversely, given a critical section $\psi\in\Gamma(M,J^{1}E^{*})$, we introduce $\phi=\sigma\circ\psi$, which is a critical section of the Lagrangian variational problem. These relationships are captured in the following commutative diagram [31]
[…]
II.4 An Example: The Free Scalar Field
The tone we have adopted thus far has been highly formal; in order to illustrate the utility of the multisymplectic formalism, and to provide a less abstract presentation of the ideas we have developed, we shall consider a simple, albeit relevant example: that of a free scalar field in a Minkowski background [57].
Spacetime coordinates on $M=\mathbb{R}^{1,3}$ are denoted $x^{\mu}$, with $\mu=0,1,2,3$; since we have a single real scalar field, the covariant configuration space $E$ is simply $E=M\times\mathbb{R}$, and we shall take local coordinates on $J^{1}E$ to be $(x^{\mu},y,y_{\mu})$. In this way, the Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$ is given by
$L(x^{\mu},y,y_{\mu})=\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}-\frac{1}{2}m^{2}y^{2}$ (20)
in which we have adopted the mostly minus convention $\eta_{\mu\nu}=\textrm{diag}(+1,-1,-1,-1)$, more commonly employed throughout the quantum field theory literature [58, 59]. The scalar energy $E_{L}\in C^{\infty}(J^{1}E)$ for this system is simply
$E_{L}=\frac{\partial L}{\partial y_{\mu}}\,y_{\mu}-L=\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}+\frac{1}{2}m^{2}y^{2}$ (21)
The Poincaré-Cartan form is of degree 4, and has the following local coordinate expression
$\Theta_{L}=\eta^{\mu\nu}y_{\nu}\,dy\wedge d^{3}x_{\mu}-\left(\frac{1}{2}\eta^{\mu\nu}y_{\mu}y_{\nu}+\frac{1}{2}m^{2}y^{2}\right)d^{4}x$ (22)
The equations of motion for a holonomic section $\phi(x)=(x^{\mu},y(x))$, calculated according to (8), read
$\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}y-m^{2}y=0$
which, as anticipated, is simply the Klein-Gordon equation for a single scalar field in flat spacetime. We also highlight that the derivatives in (8) are to be evaluated along the canonical lifting of $\phi$, which implies that $y_{\mu}\rightarrow\partial_{\mu}y$.
The momenta $p^{\mu}$ canonically conjugate to the field variable $y$ are given by
$p^{\mu}=\frac{\partial L}{\partial y_{\mu}}=\eta^{\mu\nu}y_{\nu}$
Consequently, the Hamiltonian function reads
$H=\frac{1}{2}\eta_{\mu\nu}p^{\mu}p^{\nu}+\frac{1}{2}m^{2}y^{2}$
Consider a locally decomposable, integrable multivector field $\boldsymbol{X}\in\mathfrak{X}^{4}(J^{1}E^{*})$, expressed in local coordinates as
$\boldsymbol{X}=\bigwedge\limits_{\mu=0}^{3}\left(\frac{\partial}{\partial x^{\mu}}+F_{\mu}\frac{\partial}{\partial y}+G^{\nu}_{\mu}\frac{\partial}{\partial p^{\nu}}\right)$
in which, referring to (16), $\bar{\sigma}$-transversality has been enforced setting the multiplicative function $f$ to unity. Using the following expression for the Hamilton-Cartan 5-form
$\Omega_{h}=-\,dp^{\mu}\wedge dy\wedge d^{3}x_{\mu}+(\eta_{\mu\nu}p^{\mu}\,dp^{\nu}+m^{2}y\,dy)\,\wedge\,d^{4}x$
it follows that the coordinate-free equations of motion (19) become
$F_{\mu}=\frac{\partial H}{\partial p^{\mu}}=\eta_{\mu\nu}p^{\nu}\quad\quad\quad\quad G^{\mu}_{\mu}=-\,\frac{\partial H}{\partial y}=-\,m^{2}y$
Supposing that this holonomic multivector field $\boldsymbol{X}$ has integral sections of the form $\psi(x)=\left(x^{\mu},y(x),p^{\mu}(x)\right)$, we may write
$\frac{\partial y}{\partial x^{\mu}}=F_{\mu}=\eta_{\mu\nu}p^{\nu}\quad\quad\quad\quad\frac{\partial p^{\mu}}{\partial x^{\mu}}=G^{\mu}_{\mu}=-\,m^{2}y$
and these equations may easily be shown to be equivalent to the Klein-Gordon equation, as deduced from the Lagrangian formalism.
II.5 Singular Lagrangians
The free scalar field theory of the previous example was a particularly simple case of a hyperregular Lagrangian; such functions admit globally-defined multivector field solutions which are locally decomposable, $\widehat{\pi}-$transverse and have holonomic integral sections. Many of the Lagrangian functions encountered in more physically interesting contexts are not of this nature, and require the implementation of a geometric constraint algorithm, as developed in [42, 60], for example.
As we have seen, the differential form $\Theta_{L}$ associated with a singular Lagrangian is, in general, pre-multisymplectic; as such, it is not guaranteed that $\widehat{\pi}-$transverse, integrable multivector field solutions to the equation $\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$ exist. Moreover, if such objects do exist, they are, with almost complete certainty, not defined on the entirety of $J^{1}E$. In the most favourable of cases, solutions exist on some submanifold $S_{f}\subset J^{1}E$; this is the final constraint submanifold, obtained through successive iteration of the geometric constraint algorithm.
In practice, this algorithm is implemented locally, following a number of steps, in which we define spaces of successively more restricted multivector fields. In particular, we begin by introducing the space
$\textrm{ker}^{m}\,\Omega_{L}:=\{\boldsymbol{X}_{L}\in\mathfrak{X}^{m}(J^{1}E)\;|\;\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0\,\}$
consisting of all $m$-multivector fields which are solutions to the equations of motion, but are not necessarily $\widehat{\pi}-$transverse or integrable. Within this space, we then seek the subset of those multivector fields which are locally decomposable and $\widehat{\pi}-$transverse, denoting the resulting subspace $\textrm{ker}^{m}_{\widehat{\pi}}\,\Omega_{L}\subset\textrm{ker}^{m}\,\Omega_{L}$.
Following this, we further restrict $\textrm{ker}^{m}_{\widehat{\pi}}\,\Omega_{L}$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (6). In general, this step gives rise to further consistency conditions, and possibly to new constraints; in order to ensure that the dynamics of the system remain confined to the constraint submanifold, it is necessary to impose tangency conditions on each of the constraints. This is achieved by demanding that the Lie derivative of any constraint along each of the vector fields which constitute $\boldsymbol{X}_{L}$ vanish.
Finally, we perform an analysis of the integrability of our semi-holonomic solutions, in order to reach the final subspace $\textrm{ker}^{m}_{\scriptscriptstyle H}\,\Omega_{L}$, consisting of locally decomposable, $\widehat{\pi}-$transverse, holonomic multivector fields. Examples of the local implementation of this geometric algorithm may be found in [61] and [56], for example.
Throughout this analysis, we have assumed the singular Lagrangian to be almost-regular; consequently, we know that $\mathcal{I}=\mathcal{FL}(J^{1}E)$ is a closed submanifold of $J^{1}E^{*}$. In such cases, it is possible to define the Hamilton-Cartan forms $\Theta_{h}^{\scriptscriptstyle\mathcal{I}}$ and $\Omega_{h}^{\scriptscriptstyle\mathcal{I}}$ restricted to $\mathcal{I}$, and formulate the Hamilton-Jacobi principle, as above, on the (generally) pre-multisymplectic manifold $(\mathcal{I},\Omega^{\scriptscriptstyle\mathcal{I}}_{h})$. As in the Lagrangian case, the multivector field solutions to the equations of motion are not guaranteed to exist, and in the case that they do, we must apply the above algorithm in order to deduce the appropriate final constraint submanifold of $\mathcal{I}$.
@@ III Multicontact Formulation of Classical Field Theories
Over the course of the preceding sections, we have dedicated significant effort to understanding how multisymplectic geometry may be used to formulate a rigorous description of classical field theories. Since the formalism we have developed generalises the concepts encountered in standard symplectic geometry, it should be relatively clear that all field theories treated within this framework are conservative in nature.
As we shall see, the identification of a dynamical similarity, and subsequent implementation of a symmetry reduction procedure, will inevitably necessitate the introduction of a framework which accommodates systems of a frictional (or dissipative) nature [24, 62, 63, 21]. The most appropriate arena in which to study such a class of field theories is the multicontact formalism [64, 65, 66], whose construction we shall summarise somewhat more succinctly than our treatment of multisymplectic geometry.
@@ III.1 Lagrangian Formalism
Throughout this section, we shall continue to work with the fibred manifolds $\pi:E\rightarrow M$, and $\kappa:J^{1}E\rightarrow E$, with $\textrm{dim}\,M=m$ and $\textrm{dim}\,E=n+m$. The novelty of the multicontact formalism is that the Lagrangian density is now an $m$-form on the manifold
$\mathcal{S}:=J^{1}E\times_{M}\wedge^{m-1}T^{*}M\cong J^{1}E\times\mathbb{R}^{m}$ (23)
The space $\mathcal{S}$ is of dimension $2m+n+nm$, and is a bundle over both $E$, with projection $\tau:\mathcal{S}\rightarrow E$, and $M$, with $\beta=\pi\circ\tau:\mathcal{S}\rightarrow M$. We take local coordinates on $\mathcal{S}$ to be $(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})$, in which the $s^{\mu}$ span the factor of $\mathbb{R}^{m}$ in (23), and have the field-theoretic interpretation of an action density. Pulling back the volume form $\omega\in\Omega^{m}(M)$ to $\mathcal{S}$, the Lagrangian density is written locally as
@@
The Lagrangian function $L$ is referred to as regular if the Hessian matrix
$\left(\frac{\partial^{2}L}{\partial y^{a}_{\mu}\partial y^{b}_{\nu}}\right)$
is regular at all points of $\mathcal{S}$. In close analogy to the multisymplectic case, we have the Lagrangian $m$-form $\Theta_{L}$, which, in local coordinates, is given by
$\Theta_{L}=\left(ds^{\mu}-\frac{\partial L}{\partial y^{a}_{\mu}}dy^{a}\right)\,\wedge\,d^{m-1}x_{\mu}+\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{m}x$ (25)
Note that, in contrast to the case of multisymplectic Lagrangian functions, if $L$ is singular, then the associated $m$-form $\Theta_{L}$ is not strictly required to be pre-multicontact, since the corresponding Reeb distribution is not guaranteed to exist. Such pathological examples will not be of interest to us, and so we shall consider any singular Lagrangian to define a pre-multicontact system.
Given an $r$-form $\xi\in\Omega^{r}(\mathcal{S})$, we introduce the following operator
$\begin{split}\bar{d}\;:\;\Omega^{r}(\mathcal{S})\;&\longrightarrow\;\Omega^{r+1}(\mathcal{S})\\ \xi\;&\longmapsto\;\bar{d}\xi:=d\xi+\sigma_{\Theta_{L}}\wedge\,\xi\end{split}$
where $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$ is known as the dissipation form, and is expressed in local coordinates as
@@
Note that, in general, the operator $\bar{d}$ is not nilpotent, so that there exists no sensible way to formulate cohomology groups, unless $d\sigma_{\Theta_{L}}=0$, in which case, we obtain the so-called Lichnerowicz-Jacobi cohomology [67]. Our motivation for introducing this ‘twisted‘ differential operator is that the Lagrangian $(m+1)$-form associated with $\Theta_{L}$ is not (as in the multisymplectic case) defined simply as $\Omega_{L}=-\,d\Theta_{L}$; instead, we have
$\Omega_{L}=\bar{d}\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (27)
Note that this $(m+1)$-form is defined without the minus sign used in the multisymplectic case.
Within the context of multisymplectic geometry, we define a conserved quantity to be an (m1)(m-1)-form $\zeta^{\mu}\,d^{m-1}x_{\mu}$, such that, on-shell, we have $\partial_{\mu}\zeta^{\mu}=0$ [55, 68]. Action-dependent field theories, by contrast, contain dissipated quantities; in particular, when the equations of motion are satisfied, instead of a conservation law of the form $\partial_{\mu}\zeta^{\mu}=0$, we find that
$\partial_{\mu}\zeta^{\mu}=\frac{\partial L}{\partial s^{\mu}}\zeta^{\mu}$
which is captured precisely by the dissipation form (26). It is interesting to note that all quantities within an action-dependent field theory follow the same dissipation law, and that this is determined by the Lagrangian (or Hamiltonian) function. Additionally, not all dissipation laws correspond to empirically measurable losses: this fact will be of particular relevance when discussing the physical interpretation of the symmetry-reduced models presented in later sections.
Given a regular multicontact Lagrangian system $(\mathcal{S},\Theta_{L})$, the equations of motion for a holonomic section $\Psi\in\Gamma(M,\mathcal{S})$ are given by
$\Psi^{*}\Theta_{L}=0\quad\quad\quad\textrm{and}\quad\quad\quad\Psi^{*}\iota_{Z}\Omega_{L}=0\quad\textrm{for all}\;Z\in\mathfrak{X}^{\infty}(\mathcal{S})$ (28)
Since $\Psi$ is holonomic, it may be expressed locally as
@@
with which the coordinate-free equations (28) become
$\begin{split}\frac{\partial}{\partial x^{\mu}}\left(\frac{\partial L}{\partial y^{a}_{\mu}}\circ\Psi\right)=\left(\frac{\partial L}{\partial y^{a}}+\frac{\partial L}{\partial y^{a}_{\mu}}\frac{\partial L}{\partial s^{\mu}}\right)\circ\Psi\\ \frac{\partial s^{\mu}}{\partial x^{\mu}}=L\circ\Psi\end{split}$ (29)
We shall refer to these expressions as the Herglotz-Lagrange field equations [69]. As in the multisymplectic case, the dynamical evolution of our system may also be expressed in terms of multivector fields; in the current context, the equations of motion for locally decomposable $\beta$-transverse multivector fields $\boldsymbol{X}_{L}\in\mathfrak{X}^{m}(\mathcal{S})$ are
$\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Theta_{L}=0\quad\quad\quad\quad\quad\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$ (30)

### v4

@@ II.3 Hamiltonian Field Theory
When analysing systems invariant under scaling symmetries, it will be of benefit to have at our disposal both the Lagrangian and Hamiltonian descriptions. Since the contact-reduced Hamiltonian may always be obtained by performing a Legendre transform of the corresponding Lagrangian, our treatment of the Hamiltonian construction will be heavily abridged. The relevant geometrical setting is the restricted multimomentum bundle $\sigma:J^{1}E^{*}\rightarrow E$, which serves as the covariant phase space; further details on this construction may be found in [34, 35, 36]. We introduce the projection $\bar{\sigma}:=\pi\circ\sigma:J^{1}E^{*}\rightarrow M$, and write local coordinates on $J^{1}E^{*}$ as $(x^{\mu},y^{a},p^{\mu}_{a})$. The Legendre map $\mathcal{FL}:J^{1}E\rightarrow J^{1}E^{*}$ acts on these coordinates according to
@@
A Lagrangian system is said to be regular if $\mathcal{FL}$ is a local diffeomorphism, hyperregular if this diffeomorphism is global, and singular if $\mathcal{FL}$ is not a diffeomorphism [37, 38, 39]. Throughout, we assume all systems to be (hyper)regular, and will only discuss the treatment of singular Lagrangians in our concluding remarks.
Application of the Legendre map to the Cartan forms (2) gives the corresponding Hamiltonian objects
$\Theta_{H}=p^{\mu}_{a}\,dy^{a}\wedge d^{d-1}x_{\mu}-H\,d^{d}x\quad\quad\quad\quad\Omega_{H}=-\,d\Theta_{H}$ (8)
in which $H:J^{1}E^{*}\rightarrow\mathbb{R}$ denotes the Hamiltonian function. From $\Omega_{H}$, we define the following 2-forms
$\omega_{H}^{\mu}:=-\,\iota_{\partial_{d-1}}\cdots\,\iota_{\partial_{0}}(\Omega_{H}\wedge dx^{\mu})=dy^{a}\wedge\,dp^{\mu}_{a}$ (9)
which will again be used extensively in section (VI), and should be compared to (3). The equations of motion may be derived from variational arguments [33], and if a critical section $\psi$ is expressed locally as $\psi(x)=(x^{\mu},\psi^{a}(x),\psi^{\mu}_{a}(x))$, then we find that
@@
As in the Lagrangian case, these critical sections are integral to a class of $\bar{\sigma}$-transverse, multivector fields $\{\boldsymbol{X}_{H}\}$, each of which satisfies $\iota_{\boldsymbol{X}_{H}}\Omega_{H}=0$.
@@ III Multicontact Formulation of Classical Field Theories
As mentioned previously, the elimination of a scaling degree of freedom leaves a reduced space, which naturally inherits a multicontact structure [21, 22, 23]. Physically, a theory possessing an empirically inaccessible notion of global scale is dynamically equivalent to a second, simpler theory, with no such scale, but which is frictional in nature [12, 40, 41, 17]. In this section, we briefly discuss the framework required for the description of non-conservative field theories.
@@ III.1 Lagrangian Formalism
Continuing to work with the fibered manifolds $\pi:E\rightarrow M$, and $\kappa:J^{1}E\rightarrow E$, a multicontact (or Herglotz) Lagrangian density is a $d$-form on the space
$\mathcal{S}:=J^{1}E\times_{M}\wedge^{d-1}T^{*}M\cong J^{1}E\times\mathbb{R}^{d}$ (11)
The space $\mathcal{S}$ is of dimension $2d+n+nd$, and is a bundle over both $E$, with projection $\tau:\mathcal{S}\rightarrow E$, and $M$, with $\beta=\pi\circ\tau:\mathcal{S}\rightarrow M$. We take local coordinates on $\mathcal{S}$ to be $(x^{\mu},y^{a},y^{a}_{\mu},s^{\mu})$, in which the $s^{\mu}$ have the field-theoretic interpretation of an action density. Pulling back the volume form $\omega\in\Omega^{d}(M)$ to $\mathcal{S}$, the Lagrangian density is written locally as
@@
In close analogy to the multisymplectic case, we have the Lagrangian $d$-form $\Theta_{L}$, which, in local coordinates, is given by
$\Theta_{L}=\left(ds^{\mu}-\frac{\partial L}{\partial y^{a}_{\mu}}dy^{a}\right)\,\wedge\,d^{d-1}x_{\mu}+\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{d}x$ (13)
The $(d+1)$-form $\Omega_{L}$ is constructed from the dissipation form, expressed in local coordinates as
@@
We then have
$\Omega_{L}=d\Theta_{L}+\sigma_{\Theta_{L}}\wedge\,\Theta_{L}$ (15)
A holonomic section $\Psi:M\rightarrow\mathcal{S}$ may be expressed locally as
@@
Such objects satisfy the Herglotz-Lagrange field equations
$\begin{split}\frac{\partial}{\partial x^{\mu}}\left(\frac{\partial L}{\partial y^{a}_{\mu}}\circ\Psi\right)=\left(\frac{\partial L}{\partial y^{a}}+\frac{\partial L}{\partial y^{a}_{\mu}}\frac{\partial L}{\partial s^{\mu}}\right)\circ\Psi\hskip 71.13188pt\frac{\partial s^{\mu}}{\partial x^{\mu}}=L\circ\Psi\end{split}$ (16)
which should be considered the suitably modified version of (4), adapted to accommodate the non-conservative nature of action-dependent field theories. As in the multisymplectic case, the dynamical evolution of our system may also be expressed in terms of multivector fields; in the current context, the equations of motion for locally decomposable $\beta$-transverse multivector fields $\boldsymbol{X}_{L}\in\mathfrak{X}^{d}(\mathcal{S})$ read
$\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0\quad\quad\quad\quad\quad\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$ (17)